At 7am a loaded train leaves pittsburgh at 25mph. at 1100 am another train leaves the same station on a parallel track at 65mph. how far from the station will both trains be when the second train passes the first train?
Answer by richwmiller(17219)

25(x+4)=65x
25x+100=65x
100=40x
5=2x
2.5=x
they will meet at 1:30 pm
2.5*65=162.5 miles